# Non-invertible duality defects in one, two, and three dimensions via gauging spatially modulated symmetry

### v1

@@ 4.2 Two dimensions
In this subsection, we demonstrate construction of non-invertible defects in two dimensions via gauging 00-form dipole symmetry. A new feature of such defects compared with the case of the one dimension is that we have variety of hierarchical structures of dipole and global charges that we dub dipole algebra. Note that the procedure of gauging such symmetries in two dimensions were studied in a field theoretical approach [18, 53], yet, to our knowledge, the explicit procedure to perform gauging these symmetries at the lattice level has not been discussed previously.
@@ 4.2.1 Model
Accordingly, we minimally couple the matter terms to the gauged fields via
@@
$\tau^{Z}_{\mathbf{r}}\vcentcolon=$ $\tilde{\mathcal{N}}^{Z}_{x,\mathbf{r}}\tilde{\tau^{Z}}_{\mathbf{r}},$ $\sigma^{Z}_{\mathbf{r}}\vcentcolon=\tilde{\mathcal{N}}^{Z}_{y,\mathbf{r}}\tilde{\sigma}^{Z}_{\mathbf{r}},$ $\mu^{X}_{\mathbf{p}}\vcentcolon=\tilde{\mathcal{P}}^{Z}_{\mathbf{p}}\tilde{\mu}^{Z}_{\mathbf{p}},$ (88)
@@
Further, we add the following flux operators to (81) to make sure that gauge theory is dynamically trivial:
@@ 4.2.2 Toric code with dipole symmetries
For later convenience, we briefly recall the properties of the model, $\widehat{H}_{DTC}$, (94) in the limit of $h,g_{B_{x}},g_{B_{y}}\to\infty$, before delving into the construction of the non-invertible duality defect.
@@
implying that these loops also form $1$-form dipole symmetry with dual dipole algebra. Six distinct noncontractible loops defined in (96) and (98) contribute to the nontrivial GSD, which is given by $N^{6}$. We would like to emphasize that although the GSD of the toric code with dipole symmetry was identified in UV lattice and field theoretical models [54, 16, 17, 18], the discussion on the model in view of dipole algebra, especially, $1$-form dual dipole algebra (97) has not been given previously. Such a perspective will be crucial in understanding duality defects that we will turn to next.
4.2.3 Non-invertible duality defects
After review the properties of the toric code with dipole symmetry (94), we construct the duality defects in the model (92). The way we obtain these closely parallels the ones in the previous sections. After gauging the 00-form dipole symmetry, we replace a local operator $\tilde{Z}_{\mathbf{r}}$ with the Wilson operator $\mathcal{T}_{\mathbf{r}}$ which has the following form:
@@ 4.2.3 Non-invertible duality defects
$C_{1}=\frac{1}{N^{3}}\sum_{p,q,r=1}^{N}\sum_{s,t,u=1}^{N}(\zeta_{x,y}^{Z})^{p}(\zeta_{1}^{Z})^{q}(\zeta_{2}^{Z})^{r}(\eta_{x,y}^{Z})^{s}(\eta_{1}^{Z})^{t}(\eta_{2}^{Z})^{u},$ (110)
where each $1$-form operator is given in (96) and (98). See Appendix. A for more details. This is the dipole analog of the condensation defects [48]. Note that three $1$-form operators, $\zeta^{Z}_{x,y},\zeta^{Z}_{1},\zeta^{Z}_{2}$ as well as $\eta^{Z}_{x,y},\eta^{Z}_{1},\eta^{Z}_{2}$ form dual dipole algebra (97) (99).
To summarize, the operator $D$ is the desired duality defects, i.e.,
@@
$DG^{X}_{\mathbf{r}}=X_{\mathbf{r}}D,\quad D\tau^{Z}_{\mathbf{r}}=\mathcal{N}_{x,\mathbf{r}}^{Z},\quad D\sigma^{Z}_{\mathbf{r}}=\mathcal{N}_{y,\mathbf{r}}^{Z},\quad D\mu^{Z}_{\mathbf{p}}=\mathcal{P}_{xy,\mathbf{p}}^{Z},$ (111)
@@
indicating $D$ is the non-invertible duality defect. Further, (112) is the dipole analog of fusion rules studied in the $2$-category theory. The important distinction between the previous duality defects and the one in the present case is that when taking the limit $g_{x},g_{y}\to\infty$, the operator $C$ is written as
$C=\left(\sum_{\alpha,\beta,\gamma=1}^{N}Q_{2D:0}^{\alpha}Q_{2D:x}^{\beta}Q_{2D:y}^{\gamma}\right)\times\left(\frac{1}{N^{3}}\sum_{p,q,r=1}^{N}\sum_{s,t,u=1}^{N}(\zeta_{x,y}^{Z})^{p}(\zeta_{1}^{Z})^{q}(\zeta_{2}^{Z})^{r}(\eta_{x,y}^{Z})^{s}(\eta_{1}^{Z})^{t}(\eta_{2}^{Z})^{u}\right),$ (113)
@@
We can also think of a system with 00-form dipole symmetry with the hierarchical structure of global and dipole charges is inverted, 00-form dual dipole algebra and its duality defects. Since the construction of the duality defect closely parallels the one in the previous subsections, we present how to do it succinctly.
@@
$\mathcal{K}^{Z}_{x,\mathbf{r}}\vcentcolon=Z^{\dagger}_{\mathbf{r}-\frac{\mathbf{e}_{x}}{2}}Z_{\mathbf{r}+\frac{\mathbf{e}_{x}}{2}},\quad\mathcal{K}^{Z}_{y,\mathbf{r}}\vcentcolon=Z^{\dagger}_{\mathbf{r}-\frac{\mathbf{e}_{y}}{2}}Z_{\mathbf{r}+\frac{\mathbf{e}_{y}}{2}},\quad\mathcal{K}^{Z}_{xy,\mathbf{p}}\vcentcolon=Z_{\mathbf{p}+\frac{\mathbf{e}_{x}}{2}}Z^{\dagger}_{\mathbf{p}-\frac{\mathbf{e}_{x}}{2}}Z_{\mathbf{p}+\frac{\mathbf{e}_{y}}{2}}Z^{\dagger}_{\mathbf{p}-\frac{\mathbf{e}_{y}}{2}}$ (115)
@@
Note that the two operators $G^{X}_{1,\mathbf{l}_{x}}$ and $G^{X}_{2,\mathbf{l}_{y}}$ are identical to the two flux operators $B_{\mathbf{l}_{x}}$ and $B_{\mathbf{l}_{y}}$ given in (91) when replacing Pauli $Z$ operators with Pauli $X$ ones. We further add the following flux term to ensure that gauge theory is dynamically trivial:
$\begin{split}&-g_{B}\sum_{\mathbf{r}}B_{\mathbf{r}}\vcentcolon=\\ &-g_{B}\sum_{\mathbf{r}}{\tau^{X}}_{\mathbf{r}-\mathbf{e}_{x}}({\tau}^{Z\dagger}_{\mathbf{r}})^{2}{\tau^{Z}}_{\mathbf{r}+\mathbf{e}_{x}}\times{\sigma}^{Z}_{\mathbf{r}-\mathbf{e}_{y}}({\sigma}^{Z\dagger}_{\mathbf{r}})^{2}{\sigma}^{Z}_{\mathbf{r}+\mathbf{e}_{y}}\times({\mu}^{Z}_{\mathbf{p}-\frac{\mathbf{e}_{x}}{2}+\frac{\mathbf{e}_{y}}{2}})^{\dagger}{\mu}^{Z}_{\mathbf{p}-\frac{\mathbf{e}_{x}}{2}-\frac{\mathbf{e}_{y}}{2}}({\mu}^{Z}_{\mathbf{p}+\frac{\mathbf{e}_{x}}{2}-\frac{\mathbf{e}_{y}}{2}})^{\dagger}{\mu}^{Z}_{\mathbf{p}+\frac{\mathbf{e}_{x}}{2}+\frac{\mathbf{e}_{y}}{2}}.\end{split}$ (123)
@@
$\widehat{H}_{0}^{\prime}\vcentcolon=$ $-J_{x}\sum_{\mathbf{r}}\mathcal{K}^{Z}_{x,\mathbf{r}}-J_{y}\sum_{\mathbf{r}}\mathcal{K}^{Z}_{y,\mathbf{r}}-J_{xy}\sum_{\mathbf{p}}\mathcal{K}^{Z}_{xy,\mathbf{p}}-h_{x}\sum_{\mathbf{l}_{x}}X_{\mathbf{l}_{x}}-h_{y}\sum_{\mathbf{l}_{y}}X_{\mathbf{l}_{y}}-{J}_{x}\sum_{\mathbf{r}}\sigma_{\mathbf{r}}-{J}_{y}\sum_{\mathbf{r}}\tau^{Z}_{\mathbf{r}}-J_{xy}\sum_{\mathbf{p}}\mu^{Z}_{\mathbf{p}}+h.c.$
@@
Now we construct the duality defect. Similar to discussion presented in Sec. 4.2.3, we replace local operators $\tilde{Z}_{\mathbf{l}_{x}}$ and $\tilde{Z}_{\mathbf{l}_{y}}$ with the Wilson operators $W_{\mathbf{l}_{x}}$ and $W_{\mathbf{l}_{y}}$ consisting of the string of the gauge fields (see also two configurations in the bottom of Fig. 6(c)):
@@
$W_{\mathbf{l}_{y}}$ $=$ $\prod_{\hat{x}^{\prime}=0}^{\hat{x}}\left[\sigma^{Z\dagger}_{(\hat{x}^{\prime},-\frac{1}{2})}\sigma^{Z}_{(\hat{x}^{\prime},\frac{1}{2})}\right]^{\hat{x}^{\prime}}\mu^{Z\dagger}_{(\hat{x}^{\prime}+\frac{1}{2},\frac{1}{2})}\times\prod_{\hat{y}^{\prime}=1}^{\hat{y}}\tau^{Z}_{(\hat{x},\hat{y}^{\prime})}.$ (126)
@@
We check how it acts on the spins to see whether it behaves the duality defect. In some cases, such as the one when $S$ acts on spin around the boundary, it picks up non-local charges. Omitting the details of this discussion as it closely parallels the ones in the previous argument (Sec. 4.2.3), it turns out that to have the proper duality defect, one needs to multiply $S$ with some of the operators, that is,
@@
which are portrayed in Fig. 6(d). These loops are subject to the following relation:
@@
implying that the loops constitute $1$-form dipole algebra with the hierarchical structure is inverted compared with the one in the original 00-form dipole symmetry (Fig. 6(e)). Likewise, the last three loops that enter in (130) are given by
@@
The operator $D$ is the desired duality defect whose action on local spins and spin couplings reads
$DX_{\mathbf{l}_{x}}=G^{X}_{1,\mathbf{l}_{x}},\quad DX_{\mathbf{l}_{y}}=G^{X}_{2,\mathbf{l}_{y}},\quad D\mathcal{K}^{Z}_{x,\mathbf{r}}=\sigma^{Z}_{\mathbf{r}},\quad D\mathcal{K}^{Z}_{y,\mathbf{r}}=\tau^{Z}_{\mathbf{r}}D,\quad D\mathcal{K}^{Z}_{xy,\mathbf{p}}=\mu^{Z}_{\mathbf{p}}$
$DG^{X}_{\mathbf{l}_{x}}=X_{\mathbf{l}_{x}},\quad DG^{X}_{\mathbf{l}_{y}}=X_{\mathbf{l}_{y}},\quad D\sigma^{Z}_{\mathbf{r}}=\mathcal{K}^{Z}_{x,\mathbf{r}},\quad D\tau^{Z}_{\mathbf{r}}=\mathcal{K}^{Z}_{y,\mathbf{r}},\quad D\mu^{Z}_{\mathbf{p}}=\mathcal{K}^{Z}_{xy,\mathbf{p}}.$ (134)
Furthermore, this duality defect is non-invertible; one obtains the following fusion rules:

### v2

@@ 4.2 Two dimensions
In this subsection, we demonstrate construction of noninvertible operators in two dimensions via gauging 00-form dipole symmetry. A new feature of such operators compared with the case of the one dimension is that we have variety of hierarchical structures of dipole and global charges that we dub dipole algebra. Note that the procedure of gauging such symmetries in two dimensions were studied in a field theoretical approach [19, 60], yet, to our knowledge, the explicit procedure to perform gauging these symmetries at the lattice level has not been discussed previously.
@@ 4.2.1 Model
The intuition behind the Gauss law term (93) is that one decomposes charges (90) into local one with introducing extended Hilbert spaces, corresponding to gauge fields. This is in line with the fact that gauging is a procedure to promote global symmetries into local ones, and practically, on a lattice one "crops" a global spin flip into a local spin flip with inclusion of extended Hilbert spaces. Introducing
$\mathcal{G}_{2Ddip,\mathbf{r}}\vcentcolon=\tilde{\tau}^{X\dagger}_{\mathbf{r}-\mathbf{e}_{x}}(\tilde{\tau}^{X}_{\mathbf{r}})^{2}\tilde{\tau}^{X\dagger}_{\mathbf{r}+\mathbf{e}_{x}}\times\tilde{\sigma}^{X\dagger}_{\mathbf{r}-\mathbf{e}_{y}}(\tilde{\sigma}^{X}_{\mathbf{r}})^{2}\tilde{\sigma}^{X\dagger}_{\mathbf{r}+\mathbf{e}_{y}}\times\tilde{\mu}^{X}_{\mathbf{p}-\frac{\mathbf{e}_{x}}{2}+\frac{\mathbf{e}_{y}}{2}}(\tilde{\mu}^{X}_{\mathbf{p}-\frac{\mathbf{e}_{x}}{2}-\frac{\mathbf{e}_{y}}{2}})^{\dagger}\tilde{\mu}^{X}_{\mathbf{p}+\frac{\mathbf{e}_{x}}{2}-\frac{\mathbf{e}_{y}}{2}}(\tilde{\mu}^{X}_{\mathbf{p}+\frac{\mathbf{e}_{x}}{2}+\frac{\mathbf{e}_{y}}{2}})^{\dagger}\tilde{X}_{\mathbf{r}},$ (94)
one finds
$Q_{2D:0}=\prod_{\hat{x}=1}^{L_{x}}\prod_{\hat{y}=1}^{L_{y}}\mathcal{G}_{2Ddip,\mathbf{r}},\quad Q_{2D:x}=\prod_{\hat{x}=1}^{L_{x}}\prod_{\hat{y}=1}^{L_{y}}(\mathcal{G}_{2Ddip,\mathbf{r}})^{\hat{x}},\quad Q_{2D:y}=\prod_{\hat{x}=1}^{L_{x}}\prod_{\hat{y}=1}^{L_{y}}(\mathcal{G}_{2Ddip,\mathbf{r}})^{\hat{y}}.$ (95)
The Gauss law is imposed by setting $\mathcal{G}_{2Ddip,\mathbf{r}}=1\quad\forall\mathbf{r}$.
We minimally couple the matter terms to the gauged fields via
@@
$\tau^{Z}_{\mathbf{r}}\vcentcolon=$ $\tilde{\mathcal{N}}^{Z}_{x,\mathbf{r}}\tilde{\tau^{Z}}_{\mathbf{r}},$ $\sigma^{Z}_{\mathbf{r}}\vcentcolon=\tilde{\mathcal{N}}^{Z}_{y,\mathbf{r}}\tilde{\sigma}^{Z}_{\mathbf{r}},$ $\mu^{Z}_{\mathbf{p}}\vcentcolon=\tilde{\mathcal{P}}^{Z}_{\mathbf{p}}\tilde{\mu}^{Z}_{\mathbf{p}},$ (97)
@@
Further, we add the following flux operators to (88) to make sure that gauge theory is dynamically trivial (viz flatness condition of the gauge fields is imposed):
@@ 4.2.2 Toric code with dipole symmetries
For later convenience, we briefly recall the properties of the model, $\widehat{H}_{DTC}$, (103) in the limit of $h,g_{B_{x}},g_{B_{y}}\to\infty$, before delving into the construction of the noninvertible operator.
@@
implying that these loops also form $1$-form dipole symmetry with dual dipole algebra. Six distinct noncontractible loops defined in (105) and (107) contribute to the nontrivial GSD, which is given by $N^{6}$. We would like to emphasize that although the GSD of the toric code with dipole symmetry was identified in UV lattice and field theoretical models [61, 17, 18, 19], the discussion on the model in view of dipole algebra, especially, $1$-form dual dipole algebra (106) has not been given previously. Such a perspective will be crucial in understanding operators that we will turn to next.
4.2.3 noninvertible operators
After review the properties of the toric code with dipole symmetry (103), we construct the operators in the model (101). The way we obtain these closely parallels the ones in the previous sections. After gauging the 00-form dipole symmetry, we replace a local operator $\tilde{Z}_{\mathbf{r}}$ with the Wilson operator $\mathcal{T}_{\mathbf{r}}$ which has the following form:
@@
$C_{1}=\frac{1}{N^{3}}\sum_{p,q,r=1}^{N}\sum_{s,t,u=1}^{N}(\xi_{x,y}^{Z})^{p}(\xi_{1}^{Z})^{q}(\xi_{2}^{Z})^{r}(\eta_{x,y}^{Z})^{s}(\eta_{1}^{Z})^{t}(\eta_{2}^{Z})^{u},$ (119)
where each $1$-form operator is given in (105) and (107). See Appendix. A for more details. This is the dipole analog of the condensation defects [55]. Note that three $1$-form operators, $\xi^{Z}_{x,y},\xi^{Z}_{1},\xi^{Z}_{2}$ as well as $\eta^{Z}_{x,y},\eta^{Z}_{1},\eta^{Z}_{2}$ form dual dipole algebra (106) (108).
To summarize, the operator $D$ is the desired swap operator, i.e.,
@@
$DG^{X}_{\mathbf{r}}=X_{\mathbf{r}}D,\quad D\tau^{Z}_{\mathbf{r}}=\mathcal{N}_{x,\mathbf{r}}^{Z}D,\quad D\sigma^{Z}_{\mathbf{r}}=\mathcal{N}_{y,\mathbf{r}}^{Z}D,\quad D\mu^{Z}_{\mathbf{p}}=\mathcal{P}_{\mathbf{p}}^{Z}D,$ (120)
@@
indicating $D$ is the noninvertible operator. Further, (121) is the dipole analog of fusion rules studied in the $2$-category theory. The important distinction between the previous operators and the one in the present case is that when taking the limit $g_{x},g_{y}\to\infty$, the operator $C$ is written as
$C=\left(\sum_{\alpha,\beta,\gamma=1}^{N}Q_{2D:0}^{\alpha}Q_{2D:x}^{\beta}Q_{2D:y}^{\gamma}\right)\times\left(\frac{1}{N^{3}}\sum_{p,q,r=1}^{N}\sum_{s,t,u=1}^{N}(\xi_{x,y}^{Z})^{p}(\xi_{1}^{Z})^{q}(\xi_{2}^{Z})^{r}(\eta_{x,y}^{Z})^{s}(\eta_{1}^{Z})^{t}(\eta_{2}^{Z})^{u}\right),$ (122)
@@
We can also think of a system with 00-form dipole symmetry with the hierarchical structure of global and dipole charges is inverted, 00-form dual dipole algebra and its noninvertible operators. Since the construction of the operator closely parallels the one in the previous subsections, we present how to do it succinctly.
@@
$\mathcal{K}^{Z}_{x,\mathbf{r}}\vcentcolon=Z^{\dagger}_{\mathbf{r}-\frac{\mathbf{e}_{x}}{2}}Z_{\mathbf{r}+\frac{\mathbf{e}_{x}}{2}},\quad\mathcal{K}^{Z}_{y,\mathbf{r}}\vcentcolon=Z^{\dagger}_{\mathbf{r}-\frac{\mathbf{e}_{y}}{2}}Z_{\mathbf{r}+\frac{\mathbf{e}_{y}}{2}},\quad\mathcal{K}^{Z}_{xy,\mathbf{p}}\vcentcolon=Z_{\mathbf{p}-\frac{\mathbf{e}_{x}}{2}}Z^{\dagger}_{\mathbf{p}+\frac{\mathbf{e}_{x}}{2}}Z_{\mathbf{p}+\frac{\mathbf{e}_{y}}{2}}Z^{\dagger}_{\mathbf{p}-\frac{\mathbf{e}_{y}}{2}}$ (124)
@@
Note that the two operators $G^{X}_{1,\mathbf{l}_{x}}$ and $G^{X}_{2,\mathbf{l}_{y}}$ are identical to the two flux operators $B_{\mathbf{l}_{x}}$ and $B_{\mathbf{l}_{y}}$ given in (100) when replacing Pauli $Z$ operators with Pauli $X$ ones. We further add the following flux term to ensure that gauge theory is dynamically trivial, i.e., the flatness condition is imposed:
$\begin{split}&-g_{B}\sum_{\mathbf{r}}B_{\mathbf{r}}\vcentcolon=\\ &-g_{B}\sum_{\mathbf{r}}{\tau^{Z}}_{\mathbf{r}-\mathbf{e}_{x}}({\tau}^{Z\dagger}_{\mathbf{r}})^{2}{\tau^{Z}}_{\mathbf{r}+\mathbf{e}_{x}}\times{\sigma}^{Z}_{\mathbf{r}-\mathbf{e}_{y}}({\sigma}^{Z\dagger}_{\mathbf{r}})^{2}{\sigma}^{Z}_{\mathbf{r}+\mathbf{e}_{y}}\times({\mu}^{Z}_{\mathbf{p}-\frac{\mathbf{e}_{x}}{2}+\frac{\mathbf{e}_{y}}{2}})^{\dagger}{\mu}^{Z}_{\mathbf{p}-\frac{\mathbf{e}_{x}}{2}-\frac{\mathbf{e}_{y}}{2}}({\mu}^{Z}_{\mathbf{p}+\frac{\mathbf{e}_{x}}{2}-\frac{\mathbf{e}_{y}}{2}})^{\dagger}{\mu}^{Z}_{\mathbf{p}+\frac{\mathbf{e}_{x}}{2}+\frac{\mathbf{e}_{y}}{2}}.\end{split}$ (132)
@@
$\widehat{H}_{0}^{\prime}\vcentcolon=$ $-J_{x}\sum_{\mathbf{r}}\mathcal{K}^{Z}_{x,\mathbf{r}}-J_{y}\sum_{\mathbf{r}}\mathcal{K}^{Z}_{y,\mathbf{r}}-J_{xy}\sum_{\mathbf{p}}\mathcal{K}^{Z}_{xy,\mathbf{p}}-h_{x}\sum_{\mathbf{l}_{x}}X_{\mathbf{l}_{x}}-h_{y}\sum_{\mathbf{l}_{y}}X_{\mathbf{l}_{y}}-{J}_{x}\sum_{\mathbf{r}}\sigma_{\mathbf{r}}^{Z}-{J}_{y}\sum_{\mathbf{r}}\tau^{Z}_{\mathbf{r}}-J_{xy}\sum_{\mathbf{p}}\mu^{Z}_{\mathbf{p}}+h.c.$
@@
Now we construct the operator. Similar to discussion presented in Sec. 4.2.3, we replace local operators $\tilde{Z}_{\mathbf{l}_{x}}$ and $\tilde{Z}_{\mathbf{l}_{y}}$ with the Wilson operators $W_{\mathbf{l}_{x}}$ and $W_{\mathbf{l}_{y}}$ consisting of the string of the gauge fields (see also two configurations in the bottom of Fig. 6(c)) [recall that $\mathbf{l}_{x}=(\hat{x}+\frac{1}{2},\hat{y})$, and $\mathbf{l}_{y}=(\hat{x},\hat{y}+\frac{1}{2})$]:
@@
$W_{\mathbf{l}_{y}}$ $=$ $\prod_{\hat{x}^{\prime}=0}^{\hat{x}}\left[\sigma^{Z\dagger}_{(\hat{x}^{\prime},0)}\sigma^{Z}_{(\hat{x}^{\prime},1)}\right]^{\hat{x}^{\prime}}\mu^{Z\dagger}_{(\hat{x}^{\prime}+\frac{1}{2},\frac{1}{2})}\times\prod_{\hat{y}^{\prime}=1}^{\hat{y}}\tau^{Z}_{(\hat{x},\hat{y}^{\prime})}.$ (135)
@@
We check how it acts on the spins to see whether it behaves the mapping corresponding gauging. In some cases, such as the one when $S$ acts on spin around the boundary, it picks up non-local charges. Omitting the details of this discussion as it closely parallels the ones in the previous argument (Sec. 4.2.3), it turns out that to have the proper operator, one needs to multiply $S$ with some of the operators, that is,
@@
which are portrayed in Fig. 6(d). When $g_{B}\to\infty$, these loops are subject to the following relation:
@@
implying that the loops constitute $1$-form dipole algebra with the hierarchical structure is inverted compared with the one in the original 00-form dipole symmetry (Fig. 6(e)) . The last two relations in (141) are obtained by the fact that the operators are topological, i.e., independent operators depend solely on the homology class. 161616 To see this more explicitly, we derive the fourth relation in (141). We start with $T_{y}\zeta_{x}^{Z}T_{y}^{-1}=\prod_{\hat{x}=1}^{L_{x}}(\sigma^{Z\dagger}_{(\hat{x},1)})^{\hat{x}}(\sigma^{Z}_{(\hat{x},2)})^{\hat{x}}\mu^{Z\dagger}_{(\hat{x}+\frac{1}{2},\frac{3}{2})}.$ The right hand side is equivalent to $\prod_{\hat{x}=1}^{L_{x}}(\sigma^{Z\dagger}_{(\hat{x},0)})^{\hat{x}}(\sigma^{Z}_{(\hat{x},1)})^{\hat{x}}\mu^{Z\dagger}_{(\hat{x}+\frac{1}{2},\frac{1}{2})}=\zeta^{Z}_{x}$ , obtained by the fact that we focus on the ground state satisfying $B_{\mathbf{r}}\ket{\Omega}=\ket{\Omega}$ with $B_{\mathbf{r}}$ being given by (132). The last relation in (141) can be analogously derived.
Likewise, the last three loops that enter in (139) are given by
@@
The operator $D$ is the desired swap operator whose action on local spins and spin couplings reads
$DX_{\mathbf{l}_{x}}=G^{X}_{1,\mathbf{l}_{x}}D,\quad DX_{\mathbf{l}_{y}}=G^{X}_{2,\mathbf{l}_{y}}D,\quad D\mathcal{K}^{Z}_{x,\mathbf{r}}=\sigma^{Z}_{\mathbf{r}}D,\quad D\mathcal{K}^{Z}_{y,\mathbf{r}}=\tau^{Z}_{\mathbf{r}}D,\quad D\mathcal{K}^{Z}_{xy,\mathbf{p}}=\mu^{Z}_{\mathbf{p}}D$
$DG^{X}_{\mathbf{l}_{x}}=X_{\mathbf{l}_{x}}D,\quad DG^{X}_{\mathbf{l}_{y}}=X_{\mathbf{l}_{y}}D,\quad D\sigma^{Z}_{\mathbf{r}}=\mathcal{K}^{Z}_{x,\mathbf{r}}D,\quad D\tau^{Z}_{\mathbf{r}}=\mathcal{K}^{Z}_{y,\mathbf{r}}D,\quad D\mu^{Z}_{\mathbf{p}}=\mathcal{K}^{Z}_{xy,\mathbf{p}}D.$ (143)
Furthermore, this operator is noninvertible; one obtains the following fusion rules: